We are currently (2049 on 3 August) approaching station #24 on the edge of the continental shelf on transect line 5 after spending a good part of the day in the deep (3647 m) offshore waters of station #23. Our position is -66° 52.125S; -72° 35.206W. The air temperature is -1.0°C and the wind is out of the north-northwest (335) at 18 to 20 kts. The barometer has been falling a bit today and is now reading 978.7 mlb. A light snow has been coming down off and on all day and into this evening, and on the unheated decks, an inch or two has accumulated.
Yesterday (2 August) work was completed at only two stations #21 and 22 primarily of the long steams between stations #21, 22 and 23, and because of the increased time it took to do the CTD cast to the bottom at station #22 and a deep live Tucker Trawl tow (to 1000 meters). Station #21 was noteworthy because BIOMAPER-II was returned to service after 72 hours of work to repair the severe damage that occurred when the towing wire was snagged by ice filling the ship's wake and the towed body hit the bottom of the ice pack a high speed before the ship could be stopped. To improve the reaction time, an observer now is posted to watch the towing wire during the tows and to report immediately when an ice hang-up occurs to the bridge.
John Klinck reports that on 2 August, the CTD group did two stations today covering the outer shelf along the 380 line, the fourth across shelf section since sampling began. The salinometer is now working well, thanks to the efforts of Jonnette Tuft, so the salinity samples are being run. The backlog of samples should be cleared today and a report on salinity calibration should follow in a day or so.
First, some exciting news. It appears that a relatively new lump of ACC water is sitting at station 16, just off the west coast of Adelaide Island. Temperature at 300 m is about 1.88° C, which is typical of the ACC at the shelf break. Surrounding stations indicate that it is not connected to the ACC, but is isolated. This station was the place where all of the wacky vertical structure was found indicating active intrusive layering. The event log shows no net activity at this station. This is clearly a good place to survey again if there is time near the end of this cruise.
Station #20 (496 m, sampled late on 1 August) is in the middle of the shelf. The mixed layer is 80 m deep with no structure. The O2 minimum and the temperature maximum occurred at around 250 m. There are signs of layering within and just below the pycnocline.
Station #21 (476 m) in on the outer shelf and has a two part mixed layer with a total depth of about 80 m and a salinity step at about 40 m. Properties in the mixed layer on either side of this step are uniform. There is a weak O2 minimum at around 400 m, but there is no clear temperature maximum anywhere in the cast. The temperature below 300 is basically uniform at about 1.5°C. Station #22 (3329 m) is off the shelf break in the ACC. The surface mixed layer is 50 m thick with no structure. The O2 minimum and the temperature maximum occur at about 250 m. The salinity maximum occurs at 900 to 1000 m. Strong layering is observed below the mixed layer from 50 to 100 m.
Wendy Kozlowski reports that after completion of the first four grid lines, primary production measurements have been taken at a total of ten stations. Fifteen short-term, photosynthesis vs. irradiance experiments have been completed, three of which were in new, larger volume incubators, where we hope to see enough signal to read extremely low levels of production. Two 24-hour, natural light, on-deck experiments were also completed, as well as eight deployments of the Chelsea fast repetition rate fluorometer (FRRF). Sixty samples have been filtered for CHN, and with the collection and filtration assistance of the CTD group, chlorophyll samples have been filtered at all stations sampled so far.
Kendra Daly reports that on the first three survey track lines, her group participated in net and diver collections of zooplankton and completed several rate measurement experiments for the Antarctic krill, Euphausia superba. Larval E. superba were collected over much of this northern region of the study area. Abundances appear to be relatively high as was observed during the autumn GLOBEC cruises between April and June. Densities of larvae also are relatively high in association with the undersurface of the pack ice.
Since we left the study area in early June, almost all of the larval krill have developed to furcilia stage 6 (F6) and are 7 to 12 mm in length. Experimental results indicate that larval krill are continuing to molt, but they are remaining in the F6 stage (i.e., no development) and there is little or no growth. Feeding rates are also low.
Ctenophores are abundant near surface under the ice and in net tows. Several collected individuals had ingested numerous larvae, as well as copepods and fish larvae. The role of ctenophores as a predator on larval krill will continue to be investigated.
In addition to E. superba and ctenophores, net collections also contained the euphausiids E. crystallorophias and Thysanoessa macrura, copepods, amphipods, ostracods, pteropods, jelly fish, and fish larvae.
Ari Friedlaender reports that on 2 August, available daylight first allowed for observations to begin near 0900 (-66° 43.268S; -71° 44.615W) on our way to station #22. Skies began as partly cloudy with good visibility. The ice was 9/10ths concentration of small floes, cakes, and brash. The floes were not cemented or rafted together as were those we previously encountered closer to shore. It is apparent that the farther from the coast we go, the less consolidated and pressured, the pack ice becomes. The leads were more or less pockets and pools of open water, rather than the ribbon strips streaking between the vast, snow covered floes. Fog closed in our visibility to 0.5 nm at 1030. We reached station #22 at 1215 and remained there until the day's light had faded. No whales were sighted today; however, we did count one Weddell seal and 20 Adélie penguins spread out over several small groups.
Ana Sirovic reports that, in the early afternoon of 2 August, she and Jeff Otten climbed up the science mast. Jeff installed a preamplifier on the Yagi directional antenna. They were hoping that would further increase its range. She deployed only one sonobuoy, between stations 22 and 23, which was in the vicinity of one of the sea floor moorings (S5 at 66° 35.20 and 72° 42.31) in order to have simultaneous recordings. She, however, had very bad reception from the Yagi (about 3 nm) and had to switch to the Sinclair (the omnidirectional antenna) to get more range on the bouy. She monitored the sonobuoy for 1 h 10 min and heard no biological noise.
Chris Ribic and Erik Chapman report that on 2 August (JD-214), they
surveyed for 3 hours and 21 minutes between stations #21 and 22 over the
shelf break. Ice conditions varied between 8 and 10 tenths concentration.
Most of the survey was through cake ice separated by open water. Several
small groups of Adélie Penguins were observed along leads and on
floes in the open water. Again, relatively good numbers of Snow Petrels
were seen over open water in this area. More Adélie penguins and
birds were seen overall on this line compared to the previous two as we
move closer to Marguerite Bay.
|Southern Giant Petrel||2|
They surveyed for one hour last night from the bridge and saw 3 Snow
Petrels over leads.
BIOMAPER-II/MOCNESS report (P. Wiebe, C. Ashjian, and S. Gallager):
BIOMAPER-II was deployed at 0657 on 2 August at the end of the work at station #21 after a period of extensive repairs. The towyo section on the transit between station #21 was marked by four hang-ups of the towing wire on large cakes of ice, but the new procedures for watching for these events resulted in no damage to the towed body. Each time there was about 300 meters of wire out and the observer quickly assessed the situation, called to the bridge and had the ship stopped quickly. With the ship stopped to free the wire, the towed body sank rapidly to around 300 m before returning to the more normal towing depths as the ship again came up to the towing speed of around 5 kts. On the transit between stations #22 and 23, during the evening of 2 August, a new problem developed. The level wind on the winch stopped functioning just after we had deployed the fish and had about 100 meters of cable out. That meant that towyoing had to be stopped and only a horizontal tow could be done. After a long trouble shooting effort, Scott Gallager and Jay Ardi found that a junction box, which was supposed to have been water proofed, had leaked and a short had developed that caused the malfunction.
During the tows in the offshore waters, a number of krill-like patches of high volume backscattering were observed both in the mixed layer and at depths of 300 to 350 m. The near surface patches had dimension of 50 to 60 m in the vertical and 1 kilometer or so in the horizontal. The VPR images confirmed the fact that krill were in the shallow acoustic patches. In addition, the images also showed that there were substantial numbers of pteropods in the pcynocline along with numerous copepods.
MOCNESS tow #6 was conducted at station #21 early on August 2. The tow
was conducted to 333 m. Very little was caught in any of the nets. Copepods
dominated te deeper depths (333-150), with pteropods, chaetognaths, and
amphipods also observed. Pteropods were seen from 333-75 m and in the upper
25 m. Furcilia were observed in the upper 150 m, with larger krill at 75-50
m. Ctenophores were present only in the 75-150 m interval, which contrasted
with the distribution observed inshore at station #17 where ctenophores
were present throughout the water column. Small fish and small furcilia
were observed in the upper 25 m.